Costs associated with the manufacture of miniature high-sensitivity piezoresistive pressure transducers is $77,000 per year. A clever industrial engineer found that by spending $16,000 now to reconfigure the production line and reprogram two of the robotic arms, the cost will go down to $57,000 next year and $50,000 in years 2 through 5. Using an interest rate of 9% per year, determine the present worth of the savings due to the reconfiguration.
Answer #1

r = 9%

Cash outflow (present) = $ 16000

Savings in year 1 = 77000 - 57000 = $ 20000

Savings in years 2-5 = 77000 - 50000 = $ 27000

Present worth of savings = - 16000 + 20000/1.09 + 27000/1.092 + ... +20000/1.095

PW = $ 82598.57